Solve by factoring the quadratic equation:
x square -3x = 4

Answer by stanbon(75887)

Solve by factoring the quadratic equation:
x square -3x = 4
--------------------
x^2 -3x - 4 = 0
----------
x^2 -4x + x - 4 = 0
x(x-4) + (x-4) = 0
------
(x-4)(x+1) = 0
------
x = 4 or x = -1
=====================
Cheers,
Stan H.
---------------------